OpenStudy (anonymous):

plz help... 12/11=15/5x+3

OpenStudy (anonymous):

Are you trying to solve? I you are i am pretty sure i can help. at least ill try

OpenStudy (anonymous):

yes... for x

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Let's solve your equation step by step

OpenStudy (anonymous):

ok so i think u cross multiply first

OpenStudy (anonymous):

12/11 = 15/x + 3

OpenStudy (anonymous):

Step 1. Simplify bot sides of the equation

OpenStudy (anonymous):

12/11 = 3x + 3

OpenStudy (anonymous):

Step 2. Flip the equation

OpenStudy (anonymous):

3x + 3 = 12/11

OpenStudy (anonymous):

Step 3. Subtract 3 from both sides

OpenStudy (anonymous):

3x+3−3=12/11−3

OpenStudy (anonymous):

Step 4: Divide both sides by 3. 3x 3 = −21 11 3 x= −7 11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Answer: x= −7/11

OpenStudy (anonymous):

ok thank you